OpenStudy (asdiaasd):

Which of the following statements have the same result? Explain each step in solving each one. f(3) when f(x) = 2x + 2 f−1(4) when f(x) = 3x minus 4, all over 5 y + 10 = 2y + 1

OpenStudy (asdiaasd):

@3mar

OpenStudy (abb0t):

u put da 3 in da x n den u solvsiez 4 deh problum

OpenStudy (518nad):

lol

OpenStudy (abb0t):

tf r u laughing about goNAD?

OpenStudy (518nad):

my name is nadalie tyvm

OpenStudy (asdiaasd):

this isn't really helpful...

OpenStudy (abb0t):

more like gonads

OpenStudy (518nad):

@asdiaasd f(3) = 2(3) + 2 =?

OpenStudy (3mar):

Well, I am here.

OpenStudy (asdiaasd):

I need help with the second equation

OpenStudy (518nad):

okay f^-1(4) means f inverve function with input of 4

OpenStudy (518nad):

f(x)=(3x-4)/5 what is x if f(x) = 4

OpenStudy (asdiaasd):

4?

OpenStudy (518nad):

no solve for x

OpenStudy (518nad):

f(x)=(3x-4)/5 4=(3x-4)/5 20=3x-4, ill lte u continue

OpenStudy (asdiaasd):

this is what I got f^-1(4) when f(x)= 3x-4 all over 5 f(x)= 3x-4 all over 5 To find f^-1(x) just swap the variables to solve for x: y= 3x-4 all over 5 swap variables: x= 3y-4 all over 5 solve for x 5x=3y-4 5x+4= 3y y= 5x+4 all over 3 f^-1(x)= 5x+4 all over 3 5*4= 20 20+4= 24 24/3= 8

OpenStudy (518nad):

yup

OpenStudy (3mar):

You deserve it!

OpenStudy (518nad):

f(x)=(3x-4)/5 4=(3x-4)/5 20=3x-4, 24/3=x 8=x

OpenStudy (518nad):

remember what inverse functions are doingq

OpenStudy (asdiaasd):

okay I think I understand now I answered this question thanks

OpenStudy (518nad):

okay think about this statement f^-1(4) you are asking your self what is x so that y is 4 f(4) what is y when x is 4

OpenStudy (518nad):

this way u can avoid having to find inverse functions when its not necessary